OpenStudy (mirandaregina):

If angle A is 45 degrees and angle B 60 degrees, find sin(A)cos(B)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you know \(\sin(45^\circ)\)?

OpenStudy (mirandaregina):

@misty1212 0.7?

OpenStudy (misty1212):

i would use the exact answer \[\frac{\sqrt2}{2}\]

OpenStudy (mirandaregina):

@misty1212 ok, where do I go from there?

OpenStudy (misty1212):

find \(\cos(60^\circ)\) then multiply

OpenStudy (misty1212):

you got a unit circle cheat sheet?

OpenStudy (mirandaregina):

@misty1212 cos(60) = 1/2, √2/2 x 1/2 = √2/4

OpenStudy (misty1212):

yes

OpenStudy (mirandaregina):

@misty1212 ok thanks for helping!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]